OpenStudy (anonymous):

Let f(x) = 3x - 10. Find the value of f(-3). A. 21 B. -1 C. -19

OpenStudy (ayubie):

plug in 3 for x 3(3) - 10 9 - 10 -1

OpenStudy (solomonzelman):

No, plug in NEGATIVE 3

OpenStudy (anonymous):

yes, plug in -3... when f(-3) that means for all values of x, plug in -3 like so... f(-3)=3(-3) -10 = -9 - 10 = _________ :)

OpenStudy (anonymous):

is that not the value of f (3-)?

OpenStudy (ayubie):

wow sorry! misread that

OpenStudy (anonymous):

that alright so what is the correct answer

OpenStudy (anonymous):

thats*

OpenStudy (anonymous):

just put f(-3)=3(-3) -10 = -9 - 10 = -19

OpenStudy (anonymous):

thank you very much @engrbill

OpenStudy (anonymous):

pleasure @ sadaholic19